OpenStudy (anonymous):

Can someone help me with this: Simplify. Write your answers without exponents. (1/25)^3/2 = ? 64^-2/3 = ?

OpenStudy (anonymous):

First one \[(\frac{ 1 }{ 25 })^{\frac{ 3 }{ 2 }}\]

OpenStudy (dayakar):

25 = 5*5 = 5^2 

OpenStudy (anonymous):

second: \[64^{-\frac{ 2 }{ 3 }} \]

OpenStudy (dayakar):

64 =4*4*4 = 4^3

OpenStudy (dayakar):

CAN U DO REMAINING PART

OpenStudy (anonymous):

64?

OpenStudy (dayakar):

(64)^-2/3 = (4^3)^-2/3 = 4^3*-2/3 = 4^-2 = 1/4^2 = 1/16

OpenStudy (dayakar):

IS IT CLEAR

OpenStudy (anonymous):

yes thank you